OpenStudy (anonymous):

integral of e^(-4x)dx from 0 to 1

OpenStudy (anonymous):

hi jackie....:)

OpenStudy (anonymous):

heloo there

OpenStudy (anonymous):

so i know to use u sub

OpenStudy (anonymous):

yes... u-sub.. what answer did u get, or where did u get stuck?

OpenStudy (anonymous):

i got \[-1/4\int\limits_{0}^{-4}e^udu\]

OpenStudy (anonymous):

i dont remember where to go next

OpenStudy (anonymous):

everythings good, just evaluate interal of e^u from 0 to -4....

OpenStudy (anonymous):

but since its from 0 to -4 i thought there was another step

OpenStudy (anonymous):

there is nothing wrong with the limits just because the top one is less than the bottom one... you can still evaluate it...

OpenStudy (anonymous):

you dont have to negate the whole thing?

OpenStudy (anonymous):

but if you want, you can go from -4 to 0 and make that -1/4 outside the integral to 1/4... no, the integral is correct the way it is.... it is EQUIVALENT to: \[\frac{1}{4} \int_{-4}^0 e^u du \]

OpenStudy (anonymous):

would it be \[1/4\int\limits_{-4}^{0}e^udu\]

OpenStudy (anonymous):

hahaha weird

OpenStudy (anonymous):

it's a property of integrals... switch the limits and you'll get the opposite integral...

OpenStudy (anonymous):

ok, so i could do it the way it was

OpenStudy (anonymous):

\[\large \int_a^b f(x) dx =-\int_b^a f(x) dx \]

OpenStudy (anonymous):

yes....:)

OpenStudy (anonymous):

gotchaaa :)

OpenStudy (anonymous):

great.... :)